OpenStudy (anonymous):

Word problems make me crazy: If 4 is added to twice a number and this sum is multiplied by 2, the result is the same as if the number is multiplied by 3 and 4 is added to the product. What is the number? Use x for the variable.

OpenStudy (anonymous):

2(4 + 2x) = 3x + 4

OpenStudy (anonymous):

Then solve for x

OpenStudy (anonymous):

2(4 + 2x) = 4 + 3x

OpenStudy (anonymous):

when I solve for x I get -4, but when I check I get -12 = -8 what am i doing wrong?

OpenStudy (anonymous):

lets see 2(4 + 2x) = 4 + 3x 8 + 4x = 4 + 3x 8 + x = 4 x = -4 back to original, 4 is added to twice (-4) 4 - 8 = -4 then the sum is multiplied by 2 -4 x 2 = -8 4 + 3(-4) = -8

OpenStudy (anonymous):

took a minute for me to understand how it worked out, but I got it. Thank you for your help mridrik.

OpenStudy (anonymous):

OH! Almost forgot. Thank you for your help wombat.

OpenStudy (anonymous):

Haha yay im not forgottten! :D jk XD anytime!